verify Rolle's theorem for f on [-1, 1] defined by (x) =x^4 -4x^2 +7

1
Expert's answer
2021-11-01T17:44:35-0400

"f(x)=x^4-4x^2+7" is continuous on "[-1,1]" as a polynomial.


"f(x)=x^4-4x^2+7" is differentiable on "(-1,1)" as a polynomial.



"f(-1)=(-1)^4-4(-1)^2+7=4"

"f(1)=(1)^4-4(1)^2+7=4"


"f(-1)=4=f(1)"

Since the function "f(x)=x^4-4x^2+7" satisfies these conditions, then the function "f(x)=x^4-4x^2+7" satisfies the Rolle's theorem.

Then ther is the number "c" in "(-1, 1)" such that "f'(c)=0."



"f'(x)=(x^4-4x^2+7)'=4x^3-8x"

"f'(x)=0=>4x^3-8x=0"

"4x(x^2-2)=0"

"x_1=0, x_2=-\sqrt{2}, x_3=\sqrt{2}"

Since the function "f(x)=x^4-4x^2+7" is defined on "[-1, 1]," then "c=0" and

"f'(c)=f'(0)=0."